What percent more than 1/2 is 2/3?
It says to write the answer in a mixed number with percentages.

Found 2 solutions by mananth, josgarithmetic:
Answer by mananth(16946)

2/3 is greater than 1/2 by
2/3 - 1/2 = 1/6
% increase = (1/6)/(1/2) *100
= (1/6)*(2/1) *100
=100/3
= 33 (1/3) %

Answer by josgarithmetic(39617)